# Taming Scalable Visual Tokenizer for Autoregressive Image Generation

### v1

@@ Abstract
Figure 2: Effects of Distribution Gap on Codebook Usage. (a) T-SNE of the codebook (16,384 codebook size and 256 dimension) and sampled encoder features. (b) Codebook usage curve. The partial-update strategy adopted by VQGAN broadens the distribution gap between encoded features and non-activated codes, while those of IBQ based on all-codes updating are evenly mixed, maintaining a high codebook usage ($\sim$96%) throughout the training. Fully leveraging the codebook significantly improves the reconstruction quality.
Existing vector quantization (VQ) methods struggle with scalability, largely attributed to the instability of the codebook that undergoes partial updates during training. The codebook is prone to collapse as utilization decreases, due to the progressively widening distribution gap between non-activated codes and visual features. To solve the problem, we propose Index Backpropagation Quantization (IBQ), a new VQ method for the joint optimization of all codebook embeddings and the visual encoder. Applying a straight-through estimator on the one-hot categorical distribution between the encoded feature and codebook, all codes are differentiable and maintain a consistent latent space with the visual encoder. IBQ enables scalable training of visual tokenizers and, for the first time, achieves a large-scale codebook ($2^{18}$) with high dimension ($256$) and high utilization. Experiments on the standard ImageNet benchmark demonstrate the scalability and superiority of IBQ, achieving competitive results on both reconstruction ($1.00$ rFID) and autoregressive visual generation ($2.05$ gFID). The code and models are available at https://github.com/TencentARC/SEED-Voken.
@@ 1 Introduction
Given the remarkable scalability and generalizability of autoregressive transformers [29, 2] in large language models [17, 26], recent efforts have been made to extend this success to visual generation [28, 6, 24, 25]. The common practice is to discretize images into tokens using vector quantization (VQ) before applying autoregressive transformers for discrete sequence modeling through next-token prediction.
@@
Existing studies on autoregressive visual generation primarily aim to improve the reconstruction performance of visual tokenizers [6, 31, 32] or to refine the autoregressive modeling of transformers [25, 24, 16, 14]. The improvement of visual tokenizers is crucial, given that it can markedly amplify the potential performance of ensuing autoregressive models. Intuitively, scaling visual tokenizers by increasing the codebook size and embedding dimension could help bridge the gap between discrete and continuous representations, thereby mitigating the information loss associated with discrete tokens. However, it is noteworthy that current visual tokenizers [31, 24] have not demonstrated such scaling properties.
@@
Empirical research [6, 24] has revealed that current VQ methods struggle with scalability due to the inherent tendency of the codebook to collapse. This arises because these methods only optimize limited selected codes during each backpropagation. Such a widely-adopted partial update strategy gradually broadens the distribution gap between non-activated codes and the visual encoder’s representation space, making the non-activated codes further less likely to be selected. As shown in Fig. 2, VQGAN [6] almost fails when scaling both codebook size (i.e., $16,384$) and embedding dimension (i.e., $256$) simultaneously. Only a small amount of codes share the same distribution with the visual encoder, and the codebook usage degrades from $68\%$ to $0.002\%$ after training one epoch.
@@
To tackle the challenge, we introduce a new VQ method, namely, Index Backpropagation Quantization (IBQ). It globally updates the entire codebook in each backward process to ensure consistency with the distribution of the visual encoder. In such a way, all codes have the same probability of being selected, resulting in a high utilization of the codebook throughout the training process. Specifically, rather than directly applying the straight-through estimator [1] to the selected codes, we employ this reparameterization approach on the categorical distribution between visual features and all codebook embeddings, thereby rendering all codes differentiable. As shown in Fig. 2, the sampled visual features and the codebook embeddings from IBQ are evenly mixed. IBQ keeps a high codebook usage ($\sim 96\%$) throughout the training process. Fully leveraging the codebook effectively enhances the representation capacity, as demonstrated by the superior reconstruction of IBQ ($1.37$ rFID) compared to VQGAN ($3.98$ rFID).
Experiments on the standard ImageNet benchmark validate the scalability and superiority of our method in both reconstruction and generation. We conducted an in-depth study on the scaling law of the IBQ tokenizer from three aspects: codebook size, code dimension, and model size. We observed that the reconstruction performance or codebook usage significantly increases when scaling up tokenizer training. To our knowledge, IBQ is the pioneering work to train an extremely large codebook (i.e., $262,144$) with a relatively large code dimension (i.e., $256$). This achievement has resulted in state-of-the-art reconstruction performance, achieving a $1.00$ rFID. We employed the IBQ tokenizer in a vanilla autoregressive transformer for visual generation, training models of varying sizes from 300M to 2.1B, and achieved competitive results.
@@
  • We present a family of vanilla autoregressive image generation models ranging from 300M to 2.1B, remarkably outperforming competing methods, e.g., LlamaGen [24], and Open-MAGVIT2 [16].
@@ 2.1 Vector Quantization
At the core of visual tokenizers is vector quantization, which maps the visual signals into discrete tokens. VQ-VAE [28] proposes an encoder-quantizer-decoder structure and uses a learnable codebook as the discrete representation space. VQ-VAE2 [19] introduces multi-scale hierarchical VQ-VAE to enhance local features. VQGAN [6] further uses adversarial loss and perceptual loss for good perceptual quality. RQ-VAE [13] and DQ-VAE [9] improve VQGAN by residual quantization and dynamic quantization respectively. To improve codebook utilization for large-size codebooks, some works try to decrease code dimension [31, 24]. Following this observation, MAGVIT-v2 [32] reduces the code dimension to zero, and significantly enlarges the codebook size to $2^{18}$ with Lookup-Free Quantization. Instead of joint optimization of the model and codebook, VQGAN-LC [36] extends the codebook size to 100,000 by utilizing a frozen codebook with a trainable projector.
Existing VQ methods suffer from codebook collapse when scaling up tokenzier, and usually use a small codebook size or low code dimension, which limits the representational capacity of the codebook. In contrast, our proposed IBQ shows consistent improvements when scaling up code dimension, codebook size and model size.
@@ 2.2 Tokenized Visual Generation
Tokenizers map continuous visual signals into a discrete token sequence. For subsequent visual generation, there are two approaches, including non-autoregressive (NAR) and autoregressive (AR) generation. NAR [3, 32] usually adopts BERT-style transformers to predict masked tokens. For inference, these methods generate all tokens of an image simultaneously, and iteratively refine the generated images conditioned on the previous generation. In contrast, AR models perform next-token prediction in a raster-scan manner. VQGAN [6] adopts GPT2-medium architecture, while LlamaGen [24] employs Llama [26] for scalable image generation. VAR [25] extends “next-token prediction” to “next-scale prediction” and introduces adaptive normalization (AdaLN [18]) to improve visual generation quality. Open-MAGVIT2 [16] introduces asymmetric token factorization for super-large codebook learning. In this paper, we follow vanilla autoregressive visual generation, and adopt Llama [26] with AdaLn [18] for image generation.
@@
Figure 3: Gradient flow of different VQ methods. Naïve VQ and VQGAN/VQVAE only update the selected codes in each backward process. IBQ updates all codes simultaneously by transferring the gradients of soft one-hot categorical distribution to hard one-hot index.
@@ 3 Method
Our framework consists of two stages. The first stage is to learn a scalable visual tokenizer with high codebook utilization via Index Backpropagation Quantization. In the second stage, we use an autoregressive transformer for visual generation by next-token prediction.
3.1 Preliminary
Vector quantization (VQ) maps continuous visual signals into discrete tokens with a fixed-size codebook $\mathcal{C}\in\mathbb{R}^{K\times D}$, where $K$ is the codebook size and $D$ is the code dimension. Given an image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, VQ first utilizes an encoder to project the image into the feature map $\mathcal{Z}\in\mathbb{R}^{h\times w\times D}$, where $h=H/p$, $w=W/p$, and $p$ is the downsample ratio. Then the feature map is quantized into $\mathcal{Q}\in\mathbb{R}^{h\times w\times D}$ as discrete representations with the codebook. Finally, the decoder reconstructs the image given the quantized features.
For quantization, previous methods typically calculate the Euclidean distance between each visual feature $z\in\mathbb{R}^{D}$ and all codes of the codebook, and then select the closest code as its discrete representation [28, 6]. Since the $\operatorname*{arg\,min}$ operation in quantization is non-differentiable, they apply the straight-through estimator on selected codes to copy the gradients from the decoder, to optimize the encoder simultaneously. The quantization process is as follows:
@@ 3.1 Preliminary
The partial updating strategy (i.e., only selected codes are optimized) adopted by these methods progressively widens the distribution gap between visual features and non-activated codes. It incurs the instability during training due to the codebook collapse, which hampers the scalability of the visual tokenizer.
@@ 3.2 Index Backpropagation Quantization
Quantization. To ensure the consistent distribution between the codebook and encoded features through the training, we introduce an all-codes updating method, Index Backpropagation Quantization (IBQ). At the core of IBQ is passing gradients backward to all codes of the codebook, instead of the selected codes only. Algorithm 1 provides the pseudo-code of IBQ.
@@
Then we copy the gradients of soft one-hot categorical distribution to hard one-hot index:
@@
Given the index, the quantized feature is obtained by:
@@
where $\mathcal{L}_{R}$ is reconstruction loss of image pixels, $\mathcal{L}_{Q}$ is quantization loss between the selected code embeddings and encoded features, $\mathcal{L}_{P}$ is perceptual loss from LPIPS [34], $\mathcal{L}_{G}$ is adversarial loss with PatchGAN discriminator [10] to enhance the image quality, and $\mathcal{L}_{E}$ is entropy penalty to encourage codebook utilization.
@@
${z_{q}}^{\prime}=$ $\text{Ind}_{\text{hard}}^{T}\cdot[\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{K}],$ (9)
@@
Discussion with other VQ methods. Vector quantization (VQ) discretizes a continuous visual feature into a quantized code with a fix-size codebook. Naive VQ selects the code with the closest Euclidean distance to the given visual feature. Since the $\operatorname*{arg\,min}$ operation is non-differentiable, the gradient flow back to the encoder is truncated, and only the parameters of the decoder and the selected codes can be updated. Inspired by straight-through gradient estimator [1], VQ-VAE and VQGAN copy the gradients from the decoder to the encoder for gradient approximation. They also adopt quantization loss to optimize the selected codes towards the encoded features. Consequently, the VQ model can be trained in an end-to-end manner. Despite progress has been made, these VQ methods struggle to scale up tokenizers for better representation capacity.
As shown in Fig. 3, existing VQ methods simply optimize a limited number of codes towards encoded features within each backward process. It progressively broadens the distribution gap between the non-activated codes and the encoded features, eventually leading to codebook collapse. And this situation becomes more severe with the code dimension and codebook size increase. Rather than directly applying the straight-through estimator [1] on the selected codes, we employ this parameterization approach on the categorical distribution between visual features and all codebook embeddings to enable gradients backward to all codes. In such a way, the distribution between the whole codebook and the encoded features remains consistent throughout the training process. Consequently, IBQ achieves an extremely large codebook with high code dimension and utilization.
3.3 Autoregressive Transformer
After tokenization, the visual feature is quantized into discrete representations which are subsequently flattened in a raster-scan manner for visual generation. Given the discrete token index sequence $\mathcal{X}=\{x_{i}\}_{i=1}^{T}$, where $T=h^{\prime}\times w^{\prime}$, we leverage an autoregressive transformer to model the sequence dependency via next-token prediction. Specifically, the optimization process is to maximize the log-likelihood:
@@ 3.3 Autoregressive Transformer
Note that, since our focus is on the visual tokenizer, we adopt the vanilla architecture of autoregressive transformers akin to Llama [26] with AdaLN [18] for visual generation.
@@
Table 1: Reconstruction performance of different tokenizers on $256\times 256$ ImageNet 50k validation set. reproduced VQGAN.
@@
1,024 2.24 0.2580 99%
8,192 1.87 0.2437 98%
16,384 1.37 0.2235 96%
262,122 1.00 0.2030 84%
Table 4: Impact of codebook size.
@@ 4 Experiment
Figure 5: Scaling up visual tokenizers (e.g., codebook size, code dimension and model size) improves visual soundness of reconstruction. Zoom in for a better view.
@@
Table 7: Class-conditional generation on $256\times 256$ ImageNet. * specifies the generated images are $384\times 384$ and are resized to 256×256 for evaluation. The evaluation protocol and implementation are the same with ADM.

### v2

# Scalable Image Tokenization with Index Backpropagation Quantization

@@ Abstract
Figure 2: Effects of Distribution Gap on Codebook Usage. (a) T-SNE of the codebook (16,384 codebook size and 256 dimension) and sampled encoder features. (b) Codebook usage curve. The partial-update strategy adopted by VQGAN broadens the distribution gap between encoder features and non-activated codes, while those of IBQ based on all-codes updating are evenly mixed, maintaining a high codebook usage ($\sim$96%) throughout the training. Fully leveraging the codebook significantly improves the reconstruction quality.
Existing vector quantization (VQ) methods struggle with scalability, largely attributed to the instability of the codebook that undergoes partial updates during training. The codebook is prone to collapse as utilization decreases, due to the progressively widening distribution gap between non-activated codes and visual features. To solve the problem, we propose Index Backpropagation Quantization (IBQ), a new VQ method for the joint optimization of all codebook embeddings and the visual encoder. Applying a straight-through estimator on the one-hot categorical distribution between the encoded feature and codebook, all codes are differentiable and maintain a consistent latent space with the visual encoder. IBQ enables scalable training of visual tokenizers and, for the first time, achieves a large-scale codebook ($2^{18}$) with high dimension ($256$) and high utilization. Experiments on the standard ImageNet benchmark demonstrate the scalability and superiority of IBQ, achieving competitive results on reconstruction and the application of autoregressive visual generation. The code and models are available at https://github.com/TencentARC/SEED-Voken.
@@ 1 Introduction
Discrete tokenizer plays a pivotal role in processing complex data across various modalities, such as text [6, 26, 40], images [42, 11, 3], and audio [2, 5]. By transforming raw data into discrete tokens, models can effectively handle diverse data types within a unified framework [43, 6], simplifying the integration of multimodal information and facilitating native large multimodal models [38, 44].
@@
In the image domain, pioneering works like VQGAN [11] employ vector quantization (VQ) to learn visual tokenizers, enabling effective data compression and reconstruction. VQ tokenizers are deemed as the key component in applications such as autoregressive image generation [42, 11, 37, 39] and representation learning [3, 22]. However, a notable challenge with VQ-based methods is the information loss during quantization, leading to inferior reconstruction performance compared to continuous representation models like Variational Autoencoders (VAEs) [17, 31]. Intuitively, scaling visual tokenizers by increasing the codebook size and embedding dimension could help mitigate the information loss associated with discrete tokens, thereby bridging the gap between discrete and continuous representations. However, it is noteworthy that current visual tokenizers [46, 37] have not demonstrated such scaling properties.
@@
Empirical research [11, 37] has revealed that current VQ methods struggle with scalability due to the inherent tendency of the codebook to collapse. This arises because these methods only optimize a limited number of the selected codes during each backpropagation. Such a widely-adopted partial update strategy gradually broadens the distribution gap between non-activated codes and the visual encoder’s representation space, making the non-activated codes further less likely to be selected. As shown in Fig. 2, VQGAN [11] almost fails when scaling both codebook size (i.e., $16,384$) and embedding dimension (i.e., $256$) simultaneously. Only a small amount of codes share the same distribution with the visual encoder, and the codebook usage degrades from $68\%$ to $0.002\%$ after training one epoch.
@@
To tackle the challenge, we introduce a new VQ method, namely, Index Backpropagation Quantization (IBQ). It globally updates the entire codebook in each backward process to ensure consistency with the distribution of the visual encoder. In such a way, all codes have the same probability of being selected, resulting in a high utilization of the codebook throughout the training process. Specifically, rather than directly applying the straight-through estimator [4] to the selected codes, we employ this reparameterization approach on the categorical distribution between visual features and all codebook embeddings, thereby rendering all codes differentiable. As shown in Fig. 2, the sampled visual features and the codebook embeddings from IBQ are evenly mixed. IBQ keeps a high codebook usage ($\sim 96\%$) throughout the training process. Fully utilizing the codebook effectively enhances the representation capacity, as demonstrated by the superior reconstruction of IBQ ($1.37$ rFID) compared to VQGAN ($3.98$ rFID).
We conduct a comprehensive study on the scaling behavior of IBQ tokenizers along three axes: codebook size, code dimension, and model size. We observe significant gains in reconstruction quality or codebook usage when scaling up tokenizers. To our knowledge, IBQ is the pioneering work to train an extremely large codebook (i.e., $262,144$) with a relatively large code dimension (i.e., $256$). This achievement leads to state-of-the-art reconstruction quality, reaching an rFID of $1.00$. We further demonstrate the effectiveness of IBQ tokenizers in autoregressive image generation by integrating them with vanilla transformers of varying scales, ranging from 300M to 2.1B parameters, achieving competitive performance. Although IBQ is also compatible with other advanced autoregressive models, the paper does not focus on them, leaving them for future research.
@@
  • We validate the effectiveness of IBQ tokenizers in visual generation by equipping them with vanilla autoregressive models ranging from 300M to 2.1B, remarkably outperforming competing methods, e.g., LlamaGen [37], and Open-MAGVIT2 [25].
@@ 2.1 Vector Quantization
At the core of visual tokenizers is vector quantization, which maps the visual signals into discrete tokens. VQ-VAE [42] proposes an encoder-quantizer-decoder structure with a learnable codebook as the discrete representation space. VQ-VAE2 [30] introduces multi-scale hierarchical VQ-VAE to enhance local features. VQGAN [11] further uses adversarial loss and perceptual loss for good perceptual quality. RQ-VAE [21] and DQ-VAE [15] improve VQGAN by residual quantization and dynamic quantization, respectively. To improve codebook utilization for large-size codebooks, some works try to decrease code dimension [46, 37]. Following this observation, MAGVIT-v2 [47] reduces the code dimension to zero, and expands the codebook size to $2^{18}$ with Lookup-Free Quantization. Instead of joint optimization of the model and codebook, VQGAN-LC [52] extends the codebook size to 100,000 using a frozen codebook with a trainable projector. However, it introduces a bottleneck that constrains the tokenizer capacity.
Existing VQ methods suffer from codebook collapse when scaling up tokenizers and typically use small-size codebooks or low-dimensional code embeddings, limiting the representational capacity. In contrast, our proposed IBQ shows consistent improvements when scaling up codebook size, code dimension and model size.
@@ 2.2 Tokenized Visual Generation
Tokenizers map continuous visual signals into a discrete token sequence. For subsequent visual generation, there are two approaches, including non-autoregressive (NAR) and autoregressive (AR) generation. NAR [7, 47] usually adopts BERT-style transformers to predict masked tokens. For inference, these methods generate all tokens of an image simultaneously, and iteratively refine the generated images conditioned on the previous generation. In contrast, AR models perform next-token prediction in a raster-scan manner. VQGAN [11] adopts GPT2-medium architecture, while LlamaGen [37] employs Llama [40] for scalable image generation. VAR [39] extends “next-token prediction” to “next-scale prediction” and introduces adaptive normalization (AdaLN [28]) to improve generation quality. Open-MAGVIT2 [25] proposes asymmetric token factorization for super-large codebook learning.
In this paper, we adopt vanilla autoregressive models to validate the effectiveness of IBQ tokenizers in visual generation, excluding masked modeling or multi-scale structures for simplicity. Notably, IBQ is compatible with advanced generative models, which can further unlock the tokenizer potential. We leave this exploration for future work.
@@
Figure 3: Gradient flow of different VQ methods. VQGAN/VQVAE only update the selected codes in each backward process. IBQ updates all codes simultaneously by transferring the gradients of soft one-hot categorical distribution to hard one-hot index.
@@ 3 Method
3.1 Preliminary: Vector Quantization
Vector quantization (VQ) maps continuous visual signals into discrete tokens with a fixed-size codebook $\mathcal{C}\in\mathbb{R}^{K\times D}$, where $K$ is the codebook size and $D$ is the code dimension. Given an image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, VQ first utilizes an encoder to project the image into the feature map $\mathcal{Z}\in\mathbb{R}^{h\times w\times D}$, where $h=H/p$, $w=W/p$, and $p$ is the downsample ratio. The feature map is then quantized into $\mathcal{Q}\in\mathbb{R}^{h\times w\times D}$ discrete representations using the codebook. Finally, the decoder reconstructs the image given the quantized features.
Previous methods quantize each visual feature $z\in\mathbb{R}^{D}$ by selecting the nearest code from the codebook based on Euclidean distance [42, 11]. Since the $\operatorname*{arg\,min}$ operation in quantization is non-differentiable, they apply the straight-through estimator on the selected codes to copy the gradients from the decoder to the encoder, to optimize the encoder and decoder simultaneously. The quantization process can be formulated as:
@@ 3.1 Preliminary: Vector Quantization
This partial updating strategy (i.e., only selected codes are optimized) adopted by these methods progressively widens the distribution gap between visual features and non-activated codes. It incurs the instability during training due to the codebook collapse, which hampers the scalability of the visual tokenizer.
@@ 3.2 Index Backpropagation Quantization
Quantization. To ensure the consistent distribution between the codebook and encoded features through the training, we introduce an all-codes updating method, Index Backpropagation Quantization (IBQ). The core idea of IBQ is to pass gradients to all codebook embeddings, rather than the selected ones only. Algorithm 1 provides the pseudo-code of IBQ.
@@
We then copy the gradients from soft one-hot categorical distribution to hard one-hot index:
@@
Given the index, the quantized feature can be computed as:
@@
where $\mathcal{L}_{R}$ is reconstruction loss of image pixels, $\mathcal{L}_{Q}$ is quantization loss between the selected code embeddings and encoded features, $\mathcal{L}_{P}$ is perceptual loss from LPIPS [50], $\mathcal{L}_{G}$ is adversarial loss with PatchGAN discriminator [16] to enhance the image quality, and $\mathcal{L}_{E}$ is entropy penalty to encourage codebook utilization [47].
To better explain how IBQ keeps the consistent distribution between the encoder features and the whole codebook, we provide a gradient analysis. Considering the quantization loss $\mathcal{L}_{Q}=||z-z_{q}||^{2}$,
$\frac{\partial\mathcal{L}_{Q}}{\partial\mathcal{C}_{k}}$ $=-2\text{Ind}_{k}(z-z_{q})=-2p_{k}(z-z_{q}),$ (9)
$\ p_{k}$ $=\frac{\text{exp}(z^{T}\mathcal{C}_{k})}{\sum_{j=1}^{K}\text{exp}(z^{T}\mathcal{C}_{j})}.$ (10)
The softmax probabilities $p_{k}$ ensure that each $\mathcal{C}_{k}$ is updated based on its similarity to the encoder feature $z$, and $z-z_{q}$ shifts $\mathcal{C}_{k}$ toward dominant regions of the feature distribution $P_{Z}(z)$. Random batch sampling covers the whole encoder latent space, gradually aligning the entire codebook $\mathcal{C}$ with the distribution $P_{Z}(z)$ of encoder features over time.
@@
${z_{q}}^{\prime}=$ $\text{Ind}_{\text{hard}}^{T}\mathcal{C},$ (11)
@@
Discussion with other VQ methods. As shown in Fig. 3, existing VQ methods (e.g., VQ-VAE [42] and VQGAN [11]) update only a few selected codes within each backward process, progressively widening the gap between non-activated codes and encoded features, which leads to codebook collapse. This issue worsens as code dimension and codebook size increase. Instead of applying the straight-through estimator [4] on the selected codes, we employ it on the categorical distribution between visual features and all codebook embeddings to enable gradients backward to all codes. This promotes distribution consistency between the codebook and encoded features throughout training, allowing IBQ to scale up to extremely large codebook size with high code dimension and utilization.
3.3 Vanilla Autoregressive Transformer
After tokenization, the visual feature is quantized into discrete representations which are subsequently flattened in a raster-scan manner for visual generation. Given the discrete token index sequence $\mathcal{X}=\{x_{i}\}_{i=1}^{T}$, where $T=h^{\prime}\times w^{\prime}$, we employ an autoregressive transformer to model the sequence dependency through next-token prediction. Specifically, the optimization process is to maximize the log-likelihood:
@@ 3.3 Vanilla Autoregressive Transformer
Note that, since our focus is on the visual tokenizer, we adopt the vanilla architecture of autoregressive transformers akin to Llama [40] with AdaLN [28] for visual generation. More details can be deferred to the supplementary.
@@
Table 1: Reconstruction performance of different tokenizers on $\boldsymbol{256\times 256}$ ImageNet 50k validation set. reproduced VQGAN.
@@
1 1.80 0.2377 99%
2 1.55 0.2311 97%
4 1.37 0.2235 96%
Table 5: Benefit of larger model size.
@@ 4 Experiment
Figure 5: Scaling up visual tokenizers (e.g., codebook size, code dimension and model size) improves visual soundness of reconstruction.
@@
Table 7: Class-conditional generation on $\boldsymbol{256\times 256}$ ImageNet. * specifies the generated images are $384\times 384$ and are resized to 256×256 for evaluation. The evaluation protocol and implementation are the same as ADM [10].